# Robust covariance estimation and explainable outlier detection for matrix-valued data

### v1

@@ Appendix C MMCD algorithm
7:         $\pi(i)=\{\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}:i\mapsto j:d_{\pi(1)}\leq\ldots\leq d_{\pi(n)}\}$
@@
[…]
@@ Appendix D Shapley proofs
$=$ $\sum_{i=1}^{p}\sum_{j=1}^{p}\sum_{k=1}^{q}\sum_{l=1}^{q}(\hat{x}^{S\cup\{a\}}_{ik}-m_{ik})(\hat{x}^{S\cup\{a\}}_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}-\sum_{i=1}^{p}\sum_{j=1}^{p}\sum_{k=1}^{q}\sum_{l=1}^{q}(\hat{x}^{S}_{ik}-m_{ik})(\hat{x}^{S}_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
$=$ $\sum_{i\in S\cup\{a\}}\sum_{j\in S\cup\{a\}}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}-\sum_{i\in S}\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
$=$ $\sum_{i\in S\cup\{a\}}\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}+\sum_{i\in S\cup\{a\}}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{al}-m_{al})\omega_{lk}^{\operatorname{col}}\omega_{ia}^{\operatorname{row}}$
@@
$=$ $\sum_{i\in S}\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}-\sum_{i\in S}\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
$+\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ak}-m_{ak})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{aj}^{\operatorname{row}}+\sum_{i\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{al}-m_{al})\omega_{lk}^{\operatorname{col}}\omega_{ia}^{\operatorname{row}}$
@@
$=$ $2\sum_{i\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{al}-m_{al})(x_{ik}-m_{ik})\omega_{lk}^{\operatorname{col}}\omega_{ia}^{\operatorname{row}}+\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ak}-m_{ak})(x_{al}-m_{al})\omega_{lk}^{\operatorname{col}}\omega_{aa}^{\operatorname{row}}.$
@@ Appendix E Further simulation results
Simulations that analyze the behavior of affine equivariant multivariate estimators often use data generated from a standard multivariate normal distribution, i.e. mean zero and covariance identity, because the data could be transformed to this setting anyway. Lemma 3.0.1 shows that the MMCD estimators are matrix affine equivariant but we do not consider data generated from a standard matrix normal distribution. The reason for this is twofold. On the one hand, consider the fact that the ML estimators are estimated iteratively using the flip-flop algorithm (Dutilleul,, 1999), which is commonly initialized by setting either the rowwise or columnwise covariance matrix equal to the identity matrix. Hence, we would initialize one covariance matrix with the theoretically correct one, which is very unlikely in practice. On the other hand, when we know that we are dealing with data that have independent rows and/or columns, the columns and/or rows could be analyzed separately using standard multivariate techniques.
@@
This section delves deeper into the simulations outlined in Section 6. The simulations involve generating regular and outlying samples from a matrix normal distribution. A fraction, $\varepsilon$, of the clean data is replaced by outliers. The clean observations are drawn from a centered distribution, while the mean of the outliers shifts based on the parameter $\gamma$—the mean matrix of the outliers comprises entries set to $\gamma$. Three types of covariance matrices are considered: The covariance matrix $\bm{\Sigma}^{\operatorname{rnd}}$, as proposed by Agostinelli et al., (2015), is randomly generated with low correlations. The covariance matrix $\bm{\Sigma}^{\operatorname{fix}}(0.7)$ induces a relatively collinear setting, with entries defined as:
@@
While maintaining the same covariance structure for both outliers and clean samples, we explore the impact of increasing the outlier covariance by scaling the covariance of clean observations by the parameter $s$. Each simulation setting is replicated 100 times. Unless specified otherwise, we set $\bm{\Sigma}^{\operatorname{row}}=\bm{\Sigma}^{\operatorname{rnd}}$, $\bm{\Sigma}^{\operatorname{col}}=\bm{\Sigma}^{\operatorname{mix}}(0.7)$, and $s=1$ as detailed in Section 6. An overview of all parameters for the simulations is provided in Table 1. For $(p,q)=(5,20)$, all listed parameter combinations are considered, while for $(p,q)\in{(50,20),(100,50)}$, we only consider $s=1$. Additionally, for $(p,q)=(5,20)$, we compute the Minimum Covariance Determinant (MCD) estimator on the vectorized observations for comparative analysis.
@@
and angle error between eigenvalues given as
where $\hat{\bm{a}}$ and ${\bm{a}}$ are the vectors of sorted eigenvalues of $\hat{\bm{\Sigma}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}^{\operatorname{row}}$ and $\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}$, respectively. Large values of the KL divergence and the relative Frobenius error indicate difficulties in the estimation of the covariances. The angle error between the eigenvalues is in the interval $[0,1]$, and a large value means that the shape of the covariance matrix is not appropriately estimated. To assess the efficacy of outlier detection, we extend our evaluation beyond the F-score, considering precision and recall as additional metrics. Precision denotes the proportion of correctly identified outliers among all detected samples, while recall represents the proportion of correctly identified outliers among all contaminated samples.
We begin by examining the previously discussed settings in Section 6, expanding our analysis to include varying sample sizes $n\in\{20,50,100,200,300,400,500,750,1000\}$. Figure 12 presents the scores depicting covariance estimation. For the MMCD estimators the covariance estimation performance is improving with the sample size across all settings. On the other hand, the sample size has a negligible effect on the quality of the MLE estimators in the presence of outliers and a larger mean shift decreases performance. For small sample sizes, MLE and MMCD estimators are close in terms of KL divergence, but the angle error and Frobenius error indicate worse performance of the MLE estimators also for small sample sizes. The relative Frobenius error of MMCD estimators is smaller than one and thus only plotted on $[0,1]$. For the MLE estimators, it is often above one and those settings are not visible in plots.
@@
Analyzing Figure 13, we see that the recall of the MMCD estimators is close to one across all settings, except for $(p,q)=(5,20)$ and a small mean shift. The MLE estimators only detect the most severe outliers due to the masking effect, leading to a median recall below $0.25$ across all settings. With an increasing sample size, the precision of the MMCD is improving and has very low variability. On the other hand, the MLE shows very unstable results.
@@
Figure 14 shows the difference between a contamination of $\varepsilon=0.1$ and $\varepsilon=0.4$ with mean shift $\gamma=3$. The KL divergence reveals that the MMCD estimator yields more accurate results across all settings. However, for $\varepsilon=0.1$, the F-scores of the MLE are increasing with the dimensionality and perform better than the MMCD for small sample sizes. For $\varepsilon=0.4$, only the MMCD yields reliable results.
@@
For the setting with $(p,q)=(5,20)$, $\varepsilon=0.2$ and $n>100$, we also computed the MCD on the vectorized samples in addition to the matrix MLE and MMCD and considered the true mean and covariance used to generate the data as a benchmark. Figures 15 and 16 summarize the results and reveal that the MCD on the vectorized observations does not lead to robust estimators. This issue arises because the robustness of the MCD and MMCD depends on the dimensionality of the data. For the MCD it depends on $p\cdot q$ and for the MMCD it depends on $\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}$. To achieve a $99\%$ probability of obtaining at least one clean initial subset with $(p,q)=(5,20)$ and a contamination $\varepsilon=0.2$, MCD requires approximately $2.8\cdot 10^{10}$ initial subsets, while MMCD only needs 16. For the setting with the smallest mean shift, the comparison between MMCD and the actual parameters in Figure 16 highlights the difficulty of this setting, since even using the actual parameters the recall shows a lot of variability.
@@
In addition to shifting the mean of the outliers by $\gamma\in\{1,\dots,5\}$, we now consider the effect of scaling the covariance by $s\in\{1,\dots,4\}$. The difference between $s\in\{2,3,4\}$ was negligible and Figures 17 and 18 summarize the results for $s=2$. While the MLE performs quite well for outlier detection, especially compared to the setting with $s=1$ (see Figure 14), the estimated covariance matrices are not accurate. The overall performance of the MCD computed on the vectorized samples improves with increasing sample size $n$ but even more samples would be necessary to obtain similar results to the MMCD.
@@
Overall the MMCD performs very well in all settings where the outliers can be clearly distinguished from the regular observations. It could benefit from small sample corrections as in the multivariate case. Those factors could be derived in a large-scale simulation study as in Pison et al., (2002).
Figure 12: Overview of simulation results with a fraction $\varepsilon=0.2$ of contaminated samples. The quality of covariance estimation is evaluated based on the logarithm of KL divergence, angle error between eigenvalues, and relative Frobenius error.

### v2

@@ Appendix C MMCD algorithm
7:         $\pi_{1}(i)=\{\{1,\ldots,n\}\rightarrow\{1,\ldots,n\}:i\mapsto j:d_{\pi(1)}\leq\ldots\leq d_{\pi(n)}\}$
@@
7:         $(\hat{\bm{\operatorname{M}}}_{k},\hat{\bm{\Sigma}}^{\operatorname{row}}_{k},\hat{\bm{\Sigma}}^{\operatorname{col}}_{k},\bm{d}_{k},H_{k})$ = CSTEP2($\bm{\mathfrak{X}},H_{k}$) $\triangleright$ 2 MLE and C-step iterations
@@ Appendix D Shapley proofs
Proof of Proposition 5.2.1.
To show that cellwise Shapley values are not matrix affine equivariant, we consider a rowwise addition matrix $\bm{A}$ that adds the $w$-th row to the $v$-th row. For simplicity, let $\bm{B}$ be the identity matrix. Then Equation (52) yields
$((\bm{A}\bm{X})\circ(\bm{C}\bm{Y}))_{jk}=$$\begin{cases}(x_{jk}+x_{wk})(y_{jk}-y_{wk})&{j=v}\\ x_{jk}y_{jk}&{j\neq v}\end{cases}$
while
$(\bm{A}(\bm{X}\circ\bm{Y}))_{jk}$ =$=\begin{cases}x_{jk}y_{jk}+x_{wk}y_{wk}&j=v\\ x_{jk}y_{jk}&j\neq v\end{cases}.$
Hence, we do not get invariance nor equivariance for rowwise or columnwise addition matrices. This also implies that the cellwise Shapley values are not, in general, matrix affine equivariant.
Shift invariance follows from
$\bm{\Phi}(\bm{X}+\bm{C})$ $=((\bm{X}+\bm{C})-(\bm{\operatorname{M}}+\bm{C}))\circ\bm{\Omega}^{\operatorname{row}}((\bm{X}+\bm{C})-(\bm{\operatorname{M}}+\bm{C}))\bm{\Omega}^{\operatorname{col}}=\bm{\Phi}(\bm{X}),$
which means that we can assume that $\bm{X}$ has zero mean without loss of generality.
Let $\bm{Y}:=\bm{\Omega}^{\operatorname{row}}\bm{X}\bm{\Omega}^{\operatorname{col}}$, $\bm{C}:=(\bm{A}^{\prime})^{-1}$ and $\bm{D}:=(\bm{B}^{\prime})^{-1}$, then we can write the cellwise Shapley values as $\bm{\Phi}(\bm{A}\bm{X}\bm{B})=(\bm{A}\bm{X}\bm{B})\circ(\bm{C}\bm{Y}\bm{D})$. The $jk$-th entry of this matrix can be written as
$\begin{split}\phi_{jk}(\bm{A}\bm{X}\bm{B})=&((\bm{A}\bm{X}\bm{B})\circ(\bm{C}\bm{Y}\bm{D}))_{jk}=(\bm{A}\bm{X}\bm{B})_{jk}(\bm{C}\bm{Y}\bm{D})_{jk}\\ =&\sum_{i=1}^{p}\sum_{l=1}^{q}a_{ji}x_{il}b_{lk}\sum_{m=1}^{p}\sum_{n=1}^{q}c_{jm}y_{mn}d_{nk}\\ =&\sum_{i=1}^{p}\sum_{l=1}^{q}\sum_{m=1,m\neq i}^{p}\sum_{n=1,n\neq l}^{q}a_{ji}c_{jm}x_{il}y_{mn}b_{lk}d_{nk}\\ &+\sum_{i=1}^{p}\sum_{l=1}^{q}\sum_{n=1,n\neq l}^{q}a_{ji}c_{ji}x_{il}y_{in}b_{lk}d_{nk}\\ &+\sum_{i=1}^{p}\sum_{l=1}^{q}\sum_{m=1,m\neq i}^{p}a_{ji}c_{jm}x_{il}y_{ml}b_{lk}d_{lk}\\ &+\sum_{i=1}^{p}\sum_{l=1}^{q}a_{ji}c_{ji}x_{il}y_{il}b_{lk}d_{lk}.\end{split}$ (52)
If $\bm{A}$ is a scaling matrix, i.e., a diagonal matrix with non-zero entries, we have that
$a_{ji}c_{jm}=\begin{cases}1&j=i=m\\ 0&\text{otherwise}\end{cases},$
and similarly for $\bm{B}$. This implies that
$\phi_{jk}(\bm{A}\bm{X}\bm{B})=x_{jk}y_{jk}=(\bm{X}\circ\bm{Y})_{jk}=\phi_{jk}(\bm{X}),$
showing the scale invariance.
If $\bm{A}$ is a permutation matrix, i.e., a matrix consisting of any permutation of the canonical basis vectors, we have that $(\bm{A}^{\prime})^{-1}=\bm{A}$ and
$a_{ji}c_{jm}=a_{ji}a_{jm}=\begin{cases}a_{ji}&i=m\\ 0&i\neq m\end{cases},$
and similarly for $\bm{B}$. Hence Equation (52) becomes
$((\bm{A}\bm{X}\bm{B})\circ(\bm{C}\bm{Y}\bm{D}))_{jk}=\sum_{i=1}^{p}\sum_{l=1}^{q}a_{ji}c_{ji}x_{il}y_{il}b_{lk}d_{lk}=(\bm{A}(\bm{X}\circ\bm{Y})\bm{B})_{jk},$
verifying the permutation equivariance. ∎
@@
$=$ $\sum_{i=1}^{p}\sum_{j=1}^{p}\sum_{k=1}^{q}\sum_{l=1}^{q}(\hat{x}^{S\cup\{a\}}_{ik}-m_{ik})(\hat{x}^{S\cup\{a\}}_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
$-\sum_{i=1}^{p}\sum_{j=1}^{p}\sum_{k=1}^{q}\sum_{l=1}^{q}(\hat{x}^{S}_{ik}-m_{ik})(\hat{x}^{S}_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
$=$ $\sum_{i\in S\cup\{a\}}\sum_{j\in S\cup\{a\}}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
@@
$=$ $\sum_{i\in S\cup\{a\}}\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
$+\sum_{i\in S\cup\{a\}}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{al}-m_{al})\omega_{lk}^{\operatorname{col}}\omega_{ia}^{\operatorname{row}}$
$=$ $\sum_{i\in S}\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{ij}^{\operatorname{row}}$
$+\sum_{j\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ak}-m_{ak})(x_{jl}-m_{jl})\omega_{lk}^{\operatorname{col}}\omega_{aj}^{\operatorname{row}}$
$+\sum_{i\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ik}-m_{ik})(x_{al}-m_{al})\omega_{lk}^{\operatorname{col}}\omega_{ia}^{\operatorname{row}}$
@@
$=$ $2\sum_{i\in S}\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{al}-m_{al})(x_{ik}-m_{ik})\omega_{lk}^{\operatorname{col}}\omega_{ia}^{\operatorname{row}}$
$+\sum_{k=1}^{q}\sum_{l=1}^{q}(x_{ak}-m_{ak})(x_{al}-m_{al})\omega_{lk}^{\operatorname{col}}\omega_{aa}^{\operatorname{row}}.$
@@ Appendix E Further simulation results
In order to select a simulation setting, one has to consider that the ML estimators for the parameters of the matrix-variate normal distribution employ an iterative algorithm, which is commonly initialized by setting either the rowwise or columnwise covariance matrix equal to the identity matrix (Dutilleul,, 1999). Therefore, identity covariance matrices will not be used for data generation as this could lead to an undesirable advantage for the estimation.
and angle error between eigenvalues given as
where $\hat{\bm{a}}$ and ${\bm{a}}$ are the vectors of sorted eigenvalues of $\hat{\bm{\Sigma}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}^{\operatorname{row}}$ and $\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}$, respectively. Large values of the KL divergence and the relative Frobenius error indicate difficulties in the estimation of the covariances. The angle error between the eigenvalues is in the interval $[0,1]$, and a large value means that the shape of the covariance matrix is not appropriately estimated. To assess the efficacy of outlier detection, we include the F-score in addition to precision and recall. The F-score is defined as the harmonic mean of precision and recall, where precision denotes the proportion of correctly identified outliers among all detected samples, while recall represents the proportion of correctly identified outliers among all contaminated samples. The R code of the simulations and all simulation results are available in the online supplement.
E.1 Effects of dimensionality and computation time
We start by considering additional metrics for the simulations discussed in Section 6. Figure 13 shows the F-score in addition to precision and recall. The F-score shows that for $n=100$ and increasing dimensionality the robust MMCD estimators and the MLEs yield similar results. This is due to an increasing recall of the MLEs and a decreasing precision of the MMCD estimators. For $n=1000$, the F-score of the MMCD estimators is close to the benchmark and for the MCD we see the advantage of using the deterministic MCD approach over the Fast-MCD method with increasing sample size. In Figure 14 we see that the MCD performs best in all settings across all evaluation measures. For the MCD we do not see a difference in the KL divergence when swapping to the deterministic procedure. However, the angle error between eigenvalues shows clear improvements, indicating that the estimation of the shape of the covariance matrix improves. Both in terms of the angle and Frobenius error, the MCD estimator attains better scores than the MLEs even for higher $pq$, while the MLEs have better KL divergence.
We also analyze the computation times of the estimators in this setting. Figure 15 clearly shows that computation time depends on the dimensionality of the matrix-variate samples and the number of samples for all approaches. The relative increases of computation time of the matrix MLEs and the MMCD estimators are similar for $n\in\{20,100,300\}$ but for $n=1000$ the relative increase in computation time for the matrix MLEs is larger than for the MMCD estimators, highlighting the effectiveness of the subsampling approach with increasing sample size. For the MCD, we observe a decrease in computation time when $pq>300$ since the deterministic MCD is used instead of the Fast-MCD procedure. However, computing the MCD still takes longer than the MMCD approach. Hence, the matrix-variate approach does yield higher robustness and more accurate covariance estimation with shorter computation times. Although parallel processing is available for the MMCD procedure, it was not utilized in the simulations to ensure better comparability for the algorithms. Depending on the number of available threads, parallel processing yields substantial improvements in computation time.
E.2 Cellwise and block contamination
We also consider the additional metrics for the simulations comparing the three different contamination types in Figures 16 and 17. The robustness of the MMCD estimators is again confirmed using all three metrics assessing the quality of the covariance estimation. The angle error reveals that the cell contamination has less effect on the shape of the covariance matrix than the other two scenarios and that all three estimators seemingly do a good job of estimating the covariance shape. For block contamination, the MCD yields better results than the MLEs with increasing sample size and even gets close to the MMCD in terms of angle error.
Remark E.2.1.
Our cell contamination setting does not correspond to the setting of cellwise outliers (Alqallaf et al.,, 2009). We first select a subset of outlying observations and permute the cells for this selection while Alqallaf et al., (2009) select a fraction of all cells from all samples. In our setting, we can guarantee that only 10 percent of the samples are contaminated while the cellwise contamination scheme of Alqallaf et al., (2009) would likely lead to more than half of the samples being contaminated.
In further simulations, we considered different fractions of contaminated samples as well as multiple rowwise and columnwise covariance matrices for cellwise and block contamination. Additionally, we analyzed the effect of the fraction of permuted cells per observation for cell contamination, and for block contamination, we considered different mean matrices. Those simulation results are not discussed here but are available in the online supplement.
E.3 Shift outliers
@@
For shift outliers, we include an in-depth analysis of the effect of the various simulation parameters. The simulations involve generating regular and outlying samples from a matrix normal distribution. A fraction, $\varepsilon$, of the clean data is replaced by outliers. The clean observations are drawn from a centered distribution, while the mean of the outliers shifts based on the parameter $\gamma$, i.e., the mean of the outliers is set to a matrix with all entries equal to $\gamma$. Three types of covariance matrices are considered: The covariance matrix $\bm{\Sigma}^{\operatorname{rnd}}$, as proposed by Agostinelli et al., (2015), is randomly generated with low correlations. The covariance matrix $\bm{\Sigma}^{\operatorname{fix}}(0.7)$ induces a relatively collinear setting, with entries defined as:
@@
While maintaining the same covariance structure for both outliers and clean samples, we explore the impact of increasing the outlier covariance by scaling the covariance of clean observations by the parameter $s$. Each simulation setting is replicated 100 times. Unless specified otherwise, we set $\bm{\Sigma}^{\operatorname{row}}=\bm{\Sigma}^{\operatorname{rnd}}$, $\bm{\Sigma}^{\operatorname{col}}=\bm{\Sigma}^{\operatorname{mix}}(0.7)$, and $s=1$ as detailed in Section 6. An overview of all parameters for the simulations is provided in Table 1. For $(p,q)=(5,20)$, all listed parameter combinations are considered, while for $(p,q)\in{(50,20),(100,50)}$, we only consider $s=1$.
@@
We analyze the effect of the mean shift in a setting with contamination of $\varepsilon=0.2$ and compare $\gamma=1$ and $\gamma=3$. In the upper row of Figure 18, the boxplots depict F-scores across various parameter configurations. Notably, the MMCD estimators exhibit improved performance as sample sizes increase across all settings, consistently outperforming ML estimators. However, for $(p,q)=(5,20)$, in a scenario involving a minor mean shift, the F-scores derived from MMCD exhibit some volatility with larger sample sizes. This situation arises due to the proximity of outliers to regular observations, posing challenges in their identification. Notably, a more pronounced mean shift significantly simplifies outlier detection. Moreover, we see that the recall of the MMCD estimators is close to one across all settings, except for $(p,q)=(5,20)$ and a small mean shift. The MLE estimators only detect the most severe outliers due to the masking effect, leading to a median recall below $0.25$ across all settings. With an increasing sample size, the precision of the MMCD is improving and has very low variability. On the other hand, the MLE shows very unstable results.
Figure 19 presents the scores depicting covariance estimation. For the MMCD estimators the covariance estimation performance is improving with the sample size across all settings. On the other hand, the sample size has a negligible effect on the quality of the MLE estimators in the presence of outliers and a larger mean shift decreases performance. For small sample sizes, MLE and MMCD estimators are close in terms of KL divergence, but the angle error and Frobenius error indicate worse performance of the MLE estimators also for small sample sizes. The relative Frobenius error of MMCD estimators is smaller than one and thus only plotted on $[0,1]$. For the MLE estimators, it is often above one and those settings are not visible in plots.
@@
Figure 20 shows the difference between a contamination of $\varepsilon=0.1$ and $\varepsilon=0.4$ with mean shift $\gamma=1$. The KL divergence reveals that the MMCD estimator yields more accurate results across all settings. However, for $\varepsilon=0.1$, the F-scores of the MLE are increasing with the dimensionality and perform better than the MMCD for small sample sizes. For $\varepsilon=0.4$, only the MMCD yields reliable results.
@@
For the setting with $(p,q)=(5,20)$ and $\varepsilon=0.2$, we also computed the MCD on the vectorized samples in addition to the matrix MLE and MMCD and considered the true mean and covariance used to generate the data as a benchmark. Figures 21 and 22 summarize the results and reveal that the MCD on the vectorized observations does not lead to robust estimators. This issue arises because the robustness of the MCD and MMCD depends on the dimensionality of the data. For the MCD it depends on $p\cdot q$ and for the MMCD it depends on $\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}$. To achieve a $99\%$ probability of obtaining at least one clean initial subset with $(p,q)=(5,20)$ and a contamination $\varepsilon=0.2$, MCD requires approximately $2.8\cdot 10^{10}$ initial subsets, while MMCD only needs 16. For the setting with the smallest mean shift, the comparison between MMCD and the actual parameters in Figure 22 highlights the difficulty of this setting since even using the actual parameters; the recall shows a lot of variability.
@@
In addition to shifting the mean of the outliers by $\gamma\in\{1,\dots,5\}$, we now consider the effect of scaling the covariance by $s\in\{1,\dots,4\}$. The difference between $s\in\{2,3,4\}$ was negligible and Figures 23 and 24 summarize the results for $s=2$. While the MLE performs quite well for outlier detection, especially compared to the setting with $s=1$ (see Figure 20), the estimated covariance matrices are not accurate. The overall performance of the MCD computed on the vectorized samples improves with increasing sample size $n$ but even more samples would be necessary to obtain similar results to the MMCD.
@@
E.3.1 Effects of fine-grained mean shifts
To get a more in-depth view of the effect of the mean shift we consider a finer grid for the parameter $\gamma\in\{0.1,0.2,\dots,2\}$ for $n\in\{20,100,1000\}$, $(p,q)=(5,20)$, $\varepsilon=0.1$. In Figure 27, we see that for $n=20$, the MMCD has a low precision but an even higher recall than we can achieve using the actual parameters used to generate the data to compute the Mahalanobis distances for outlier detection. For larger sample sizes, the precision of the MMCD increases while the recall remains high, resulting in an F-score close to the one achieved by the actual parameters. For $n=1000$, we also computed the MCD on the vectorized observations, it attains a higher recall than the matrix MLEs but lower precision and performs worse than the MMCD in all settings. Likewise, to the results for outlier detection, Figure 28 shows similar results for covariance estimation. While the MMCD performs best in most settings it shows potential for improvement for small $n$ and $\gamma$. The simulations also show that at a level of 10 percent contamination, even a small shift $\gamma$ of the outliers negatively impacts covariance estimation and, consequently, outlier detection due to the masking effect.
E.4 Beyond normality, contaminated t-distribution
To analyze the effect of deviations from the matrix normal distribution we consider samples from a matrix t-distribution. Similar to the matrix normal distribution, the matrix t-distribution is parameterized by a mean matrix, rowwise and columnwise covariance matrices, and degrees of freedom as an additional parameter, see Gupta and Nagar, (1999) for more details. We also consider the ML estimators for the matrix t-distribution proposed by Thompson et al., (2020), which are implemented in the R package MixMatrix. We consider samples from a $p\times q=5\times 20$ centered matrix t-distribution with $\nu\in\{1,\dots,30\}$ degrees of freedom with $\bm{\Sigma}^{\operatorname{row}}=\bm{\Sigma}^{\operatorname{rnd}}\in\operatorname{PDS}(p)$ and $\bm{\Sigma}^{\operatorname{col}}=\bm{\Sigma}^{\operatorname{mix}}(0.7)\in\operatorname{PDS}(q)$ for $n\in\{20,100,1000\}$, $(p,q)=(5,20)$, $\varepsilon\in\{0.1,0.2\}$. The outliers are generated from a shifted distribution with a mean matrix of all ones with the same covariance structure and the same degrees of freedom. In Figure 29, we analyze the influence of the degrees of freedom on precision, recall, angle error between eigenvalues, and the logarithm of the relative Frobenius error for various estimators, number of samples, and levels of contamination. The angle and Frobenius error clearly show the advantage of the MMCD estimators for covariance estimation. If the distribution of the samples is known, the consistency correction outlined in Theorem 3.0.2 allows us to obtain consistency for any matrix elliptical distribution. Since we do not know the underlying distribution in practice, we use the consistency factor for the normal model given in Equation (18) which does affect the scale of the covariance but not the shape. This is also reflected in the difference between the angle and Frobenius error of the MMCD estimators and MLEs for the matrix t-distribution since the scale of the covariance has a more profound impact on the Frobenius error. In terms of angle error, the MMCD estimators perform better than the MLEs for the matrix t-distribution for all degrees of freedom $\nu$ while the MMCD shows high Frobenius errors for $\nu\leq 4$.
While the MMCD estimators and MLEs for the matrix t-distribution have a recall close to one in all settings, we see a difference in precision depending on the fraction of contaminated samples and the number of samples. For $\varepsilon=0.1$, the MLEs for the matrix t-distribution show a steep increase in precision with rising degrees of freedom for all the sample sizes. On the other hand, for $\varepsilon=0.2$, the precision is constant and low for all $n$. Similarly to the simulations based on the normal model, the precision of the MMCD estimators is low for $n=20$ and remains low for increasing degrees of freedom. While the precision increases alongside the degrees of freedom for larger sample sizes it is still low. However, this is what we would expect since the matrix t-distribution has heavier tails than the matrix normal distribution and the mean shift is rather small, such that we do not see the full potential of the MMCD estimators even under the normal model, see Section E.3.
Both the normal MLEs and the MCD estimators computed on the vectorized samples perform poorly for covariance estimation and outlier detection when the samples are generated from a matrix t-distribution.
Figure 13: Outlier detection capabilities comparing multiple matrix sizes $p\in\{2,\dots,30\}$ and $q\in\{10,20,30\}$ for $n\in\{100,1000\}$, $\gamma=1$, $\varepsilon=0.1$.
Figure 14: Quality of covariance estimation comparing multiple matrix sizes $p\in\{2,\dots,30\}$ and $q\in\{10,20,30\}$ for $n\in\{100,1000\}$, $\gamma=1$, $\varepsilon=0.1$.
Figure 15: Comparison of computation time in seconds for multiple matrix sizes $p\in\{2,\dots,30\}$ and $q\in\{10,20,30\}$ for $n\in\{20,100,300,1000\}$.
Figure 16: Quality of covariance estimation comparing block, cell, and sample contamination.
Figure 17: Outlier detection capabilities comparing block, cell, and sample contamination.